A ball is kicked by 30m/s at angle 37degree from the horizontal. find total time flight


1
Expert's answer
2021-10-24T18:24:09-0400
"t=\frac{2v\sin{37}}{g}\\t=\frac{2(30)\sin{37}}{9.8}=3.7\ s"